OpenStudy (anonymous):

Ray is 2 inch taller than Lin, and Ray is 3 inch taller than Sam. What is the average height of ray, Lin and Sam? The median height of Ray, Lin and Sam? Please help me to solve the question..

OpenStudy (anonymous):

You didn't give us enough information. We need at least 1 of their heights.

OpenStudy (anonymous):

Ray: x Lin: x-2 Sam: x-3 Average height: (x+(x-2)+(x-3))/3 = (3x-5)/3 = x-5/3 Median height: x-2

OpenStudy (anonymous):

Thank you for the quick response.. You guys have a good day..